which of the numbers 51,81,and111 is a composite number? (P.S thank you so very much)
Answer by prabhjyot(165)

A composite number is a number that has more than two factors. For example, 6 is a composite number because it has four factors (1, 2, 3, 6). 5 is not a composite number because it has only 2 factors, 1 and 5.
Add up the digits of your number, if the sum is divisible by 3, your number is composite
51 = 5+1=6
6 is divisible by 3
so 51 is composite number
81 = 8+1=9
9 is divisible by 3
so 81 is composite number
111= 1+1+1 =3
3 is divisible by 3
so 111 is composite number